# CHIQ: Contextual History Enhancement for Improving Query Rewriting in Conversational Search

### v1

@@ Abstract
In this paper, we study how open-source large language models (LLMs) can be effectively deployed for improving query rewriting in conversational search, especially for ambiguous queries. We introduce CHIQ, a two-step method that leverages the capabilities of LLMs to resolve ambiguities in the conversation history before query rewriting. This approach contrasts with prior studies that predominantly use closed-source LLMs to directly generate search queries from conversation history. We demonstrate on five well-established benchmarks that CHIQ leads to state-of-the-art results across most settings, showing highly competitive performances with systems leveraging closed-source LLMs. Our studyprovides a first step towards leveraging open-source LLMs in conversational search, as a competitive alternative to the prevailing reliance on commercial LLMs. Data, models, and source code will be publicly available upon acceptance at https://github.com/fengranMark/CHIQ.
@@ 1 Introduction
Conversational search enables users to interact with the system in a multi-turn fashion to satisfy their complex information needs Gao et al. (2022); Zamani et al. (2023). One of the crucial steps is to compose adequate search queries for each context-dependent utterance. Recent advancements in the task-solving capabilities of Large Language Models (LLMs) Ouyang et al. (2022); Shuster et al. (2022) have motivated researchers to integrate these models into existing conversational search systems.
@@
One of the main challenges in conversational search resides in the ambiguous nature of the conversation history. Figure 1 illustrates an example where solving co-reference relation in $u4$ and elaborating the response in $r3$ can help generate an adequate search query. Intuitively, performing these tasks requires basic NLP task-solving capabilities, which even small-scale open-source LLMs (e.g., 7B) possess Touvron et al. (2023); Jiang et al. (2023). The key challenge is to unlock the full capabilities of open-source LLMs for conversational search. This requires carefully preparing the conversation history to enhance its quality, rather than directly using it to generate the search query.
@@
We conduct extensive experiments using the open-source LLM, LLaMA-2-7B Touvron et al. (2023), across five well-established conversational search benchmarks under both dense and sparse retrieval settings. The experimental results indicate that enhancing the conversational history using our method achieves state-of-the-art performance across most settings, often surpassing systems powered by closed-source LLMs. Our analysis reveals that although closed-source LLMs benefit from enhancing the history, the gap with open-source models is narrower when using the enhanced history with different facets compared to the original.
@@ 2 Related Work
Different from traditional ad-hoc retrieval, which assumes users submit a stand-alone query, conversational search provides a conversational interface so that users can elaborate more complex search requirements, and interactively perform search. The main challenge lies in accurately understanding the user’s real search intent, which may be embedded within a longer, noisy, and more complex conversational context history. There are two well-established approaches in the literature for conversational search: Conversational Dense Retrieval (CDR) Qu et al. (2020); Yu et al. (2021); Mao et al. (2024) and Conversational Query Rewriting (CQR) Elgohary et al. (2019).
The CDR systems aim to fine-tune an end-to-end conversational dense retriever that can directly model the entire conversational history to return relevant documents Kim and Kim (2022); Mo et al. (2024a). Conversely, the CQR systems focus on formulating an adequate search query based on the conversational history. This query can then serve as the input to an existing, well-established retriever-ranker framework. We base our solution on CQR, leveraging its ability to integrate with existing ad-hoc search models, which has demonstrated significant practical value Dalton et al. (2022).
Earlier approaches to CQR attempted to select useful tokens from the conversation context Kumar and Callan (2020); Voskarides et al. (2020); Fang et al. (2022) or to train generative rewriter models with conversational sessions to mimic the human-rewritten query Yu et al. (2020); Lin et al. (2020); Vakulenko et al. (2021). To optimize query rewriting, some studies have adopted reinforcement learning Wu et al. (2022); Chen et al. (2022), or used the ranking signals with the rewriting model training Qian and Dou (2022); Mo et al. (2023a); Mao et al. (2023b). In addition, there have been endeavors to improve the conversion history quality through context denoising Lin et al. (2021b); Mao et al. (2022a); Krasakis et al. (2022); Mo et al. (2023b); Mao et al. (2023c), and data augmentation Dai et al. (2022); Mao et al. (2022b); Mo et al. (2024b). Unlike them, we enhance query rewriting by leveraging the NLP capabilities of open-source LLMs to reduce the ambiguity of conversational history. There have been multiple endeavors to integrate LLMs to solve traditional ad-hoc search sub-tasks Zhu et al. (2023), such as query expansion Wang et al. (2023); Gao et al. (2023), dense retrieval Ma et al. (2023); Wang et al. (2024), and re-ranking Sun et al. (2023). About conversational search, Jin et al. (2023), Jang et al. (2023), and Chen et al. (2024) attempt to improve CDR with unsupervised fine-tuning.  Mao et al. (2023a) and Ye et al. (2023) explore how LLMs can understand users’ contextualized search intents via CQR. Unlike direct rewriting the query using LLMs, we investigate various approaches for integrating refined conversational history into CQR frameworks.
@@ 3.1 Task Formulation
Let =$\mathcal{H}=\{u_{k},r_{k}\}^{n}_{k=1}$ represent the user-system conversational history, where $u_{k}$ and $r_{k}$ are the user question and the system response at the $k$-th turn. Given a new user question $q_{n+1}$, the goal of a conversational search system is to return a set of passages $\mathcal{P}_{n+1}$ that are relevant to $\mathcal{H}$ and $u_{n+1}$, which would eventually help generate the model response $r_{n+1}$. To solve the main challenge of uncovering the real search intents hidden in the user’s context-dependent query, a conversational query rewriting (CQR) module has been commonly employed as an intermediate step to obtain a rewritten query $q_{n+1}$, which in turn is used as input to an off-the-shelf retriever. Recently, LLMs have become the default option for obtaining $q_{n+1}$ as follows:
@@ 3.2.1 Question Disambiguation
Users often expect human-to-human level of interaction with modern conversational systems. They often use acronyms, ambiguous words, or coreference substitutes when asking questions, expecting native understanding and default reasoning capabilities from these systems. For a search system, the search intent is often unclear and ambiguous. Therefore, we propose a prompt to an LLM, denoted as $\mathcal{I}^{QD}$, that takes the conversational history $\mathcal{H}$ and the subsequent user question $u_{n+1}$ as input, to generate $u^{\prime}_{n+1}$, a self-contained and unambiguous version of $u_{n+1}$ which can substitute it in Eq. 1.
@@ 3.2.2 Response Expansion
In conversation sessions, it is common for model responses to be short and concise, especially for the factoid query. While brevity is often convenient for the user to acquire needed information, it makes the response less informative for a search system, which requires abundant rewrite/expansion resources. To handle this issue, we design a prompt $\mathcal{I}^{RE}$ which instructs an LLM to enrich the content of the last model response. The goal is to make it self-contained by leveraging the preceding conversational history. $\mathcal{H}^{\prime}$ can be obtained by replacing $r^{\prime}_{n}$ by $r_{n}$ in $\mathcal{H}$.
@@ 3.4 Search-Oriented Fine-tuning
Then, we select $q_{n+1}^{\prime}$, the one with the highest retrieval score determined by an off-the-shelf retriever and relevance judgment from the set of pseudo-queries $Q^{\prime}_{n+1}$ as Eq 3, which is used as the supervision signal to fine-tune a rewriting model $\mathcal{M}(\mathcal{H}^{\prime}\oplus u_{n+1})=q_{n+1}^{\prime}$ by maximum likelihood estimation. It is important to mention that the process described in this section is conducted offline and performed once, for the purpose of generating pseudo-labeled queries to fine-tune a search-oriented query rewriter. During inference, $\mathcal{H}$ and $u_{n+1}$ serve as inputs for the fine-tuned model to generate the query $q_{n+1}$, and no calls are made to the LLM, so that the latency is not much affected.
@@ 4.2 Baselines
We compared our methods with a variety of systems that can mainly be classified into three categories. More precisely, we first compare against traditional systems that fine-tune small-scale CQR models (e.g., T5-base) including: QuReTeC Voskarides et al. (2020) T5QR Lin et al. (2020), CONQRR Wu et al. (2022), ConvGQR Mo et al. (2023a), EDIRCS Mao et al. (2023b). Then, we compare with the systems that fine-tune an LLM-based retriever, e.g., create the query and document representation by the ending token from a decoder-only model, including RepLLaMA Ma et al. (2023), E5-Mistral Wang et al. (2024), and LLM-Embedder Zhang et al. (2023), or fine-tune an LLM-based CQR model as RETPO Yoon et al. (2024) and IterCQR Jang et al. (2023). Besides, we include the systems that directly obtain the rewritten query by prompting LLMs such as LLM-Aided IQR Ye et al. (2023), HyDE Gao et al. (2023), Query2doc Wang et al. (2023) and LLM4CS Mao et al. (2023a). Although not directly comparable, we report results of systems that fine-tune an ad-hoc search retriever for conversational scenarios, including the one without LLMs ConvDR Yu et al. (2021) and with LLMs InstructorR Jin et al. (2023). A detailed description of each aforementioned baseline is presented in Appendix B.3.
@@ 4.3 Implementation Details
We conduct experiments with both LLaMA-2-7B Touvron et al. (2023) and Mistral-2-7B Jiang et al. (2023) as $\mathcal{LLM(.)}$ in Eq. 1 and Eq. 2. We experiment with both BM25 Robertson et al. (2009) sparse retriever and ANCE dense retriever Xiong et al. (2020). In addition, we use FlanT5-base [^2] We report our main results using FlanT5-base to ensure the results are comparable with previous studies. Chung et al. (2022) and large models as the backbone when fine-tuning a CQR model on TopiOCQA and QReCC. The fine-tuning process consists of 10 epochs with a learning rate of 1e-5 and a batch size of 8 for both datasets. More implementation details can be found in Appendix B.2.
@@ 6 Conclusion
Potential limitations of this work include not experimenting with larger open-source LLMs, such as the 56B Mixtral Mixtral AI team (2023) or 70B LLaM a, as well as other recent models like Gemma Team et al. (2024). Additionally, the study did not incorporate more closed-source models such as GPT-4 OpenAI (2023) or Claude AnthropicAI (2023) to further study the impact of history enhancement. This is mainly due to limitations in computation (open source) and financial (close source) resources. Despite the straightforward and significant gains, some design choices could be further analyzed to potentially boost the performance even more. For instance, adding a backoff filtering strategy could detect when the LLM is producing noisy outputs, or exploring approaches that interpolate between the use of human and pseudo-queries when its higher quality as training signals for CHIQ-FT. We have considered 5 directions of enhancement in this paper. More strategies can be incorporated so that other useful enhancements can be integrated.
@@ B.3 Baselines Details
E5-Mistral Wang et al. (2024): A large ad-hoc retriever fine-tuned on top of Mistral-7B model on the synthetic dataset generated by ChatGPT-3.5 and MSMARCO.
@@ Appendix C Results of Mistral-2-7B
Table 7 and Table 8 show the performances of our methods using Mistral-2-7B as the backbone LLM, replacing LLama-2-7B as previously shown in Table 1 and Table 2, respectively. These results exhibit trends similar to those observed with LLama-2-7B, as discussed in Sections 5.1 and 5.2. More specifically, CHIQ-AD consistently outperforms CHIQ-FT across all settings, and combining the results of both methods (CHIQ-Fusion) yields the best performance. It is worth noting that results with Mistral-2-7B are systematically lower than the ones with LLama-2-7B on most datasets (except on CAsT-21). Nevertheless, performing QR on top of the enhanced history with our approach still outperforms most other settings and datasets.

### v2

@@ Abstract
In this paper, we study how open-source large language models (LLMs) can be effectively deployed for improving query rewriting in conversational search, especially for ambiguous queries. We introduce CHIQ, a two-step method that leverages the capabilities of LLMs to resolve ambiguities in the conversation history before query rewriting. This approach contrasts with prior studies that predominantly use closed-source LLMs to directly generate search queries from conversation history. We demonstrate on five well-established benchmarks that CHIQ leads to state-of-the-art results across most settings, showing highly competitive performances with systems leveraging closed-source LLMs. Our study provides a first step towards leveraging open-source LLMs in conversational search, as a competitive alternative to the prevailing reliance on commercial LLMs for query rewriting. Our code is publicly available at https://github.com/fengranMark/CHIQ.
@@ 1 Introduction
Conversational search enables users to interact with the system in a multi-turn fashion to satisfy their complex information needs Gao et al. (2022); Zamani et al. (2023). One of the crucial steps is to compose adequate search queries for each context-dependent utterance. Recent advancements in the task-solving capabilities of Large Language Models (LLMs) Ouyang et al. (2022); Chen et al. (2024b); Wang et al. (2024a); Huang et al. (2024) have motivated researchers to integrate these models into existing conversational search systems.
@@
One of the main challenges in conversational search resides in the ambiguous nature of the conversation history. Figure 1 illustrates an example where solving co-reference relation in $u_{4}$ and elaborating the response in $r_{3}$ can help generate an adequate search query. Intuitively, performing these tasks requires basic NLP task-solving capabilities, which even small-scale open-source LLMs (e.g., 7B) possess Touvron et al. (2023); Jiang et al. (2023). The key challenge is to unlock the full capabilities of open-source LLMs for conversational search. This requires carefully preparing the conversation history to enhance its quality, rather than directly using it to generate the search query.
@@
We conduct extensive experiments using the open-source LLM, LLaMA-2-7B Touvron et al. (2023), across five well-established conversational search benchmarks under both dense and sparse retrieval settings. The experimental results indicate that enhancing the conversational history using our method achieves state-of-the-art performance across most settings, often surpassing systems powered by closed-source LLMs. Our analysis reveals that although closed-source LLMs benefit from enhancing the history, the gap with open-source models is narrower when using the enhanced history with different facets compared to the original. Our contributions are summarized as follows:
  • We propose a two-step method for query rewriting that relies on open-source language models: enhancing the conversation history and then generating the search query.
  • We introduce three approaches for generating the search query on top of the enhanced conversation history: ad-hoc query rewriting (CHIQ-AD), fine-tuning a small LM for the task (CHIQ-FT), and a fusion of both approaches (CHIQ-Fusion).
  • Experiments conducted on five conversational search benchmarks demonstrate that CHIQ, using open-source LLMs, achieves state-of-the-art performance across most settings, often surpassing systems that rely on closed-source LLMs.
@@ 2 Related Work
Different from traditional ad-hoc retrieval, which assumes users submit a stand-alone query, conversational search provides a conversational interface so that users can elaborate more complex search requirements, and interactively perform search. The main challenge lies in accurately understanding the user’s real search intent, which may be embedded within a longer, noisy, and more complex conversational context history. There are two well-established approaches in the literature for conversational search: Conversational Dense Retrieval (CDR) Qu et al. (2020); Yu et al. (2021); Mao et al. (2024); Mo et al. (2024a) and Conversational Query Rewriting (CQR) Elgohary et al. (2019).
The CDR systems aim to fine-tune an end-to-end conversational dense retriever that can directly model the entire conversational history to return relevant documents Kim and Kim (2022); Mo et al. (2024b). Conversely, the CQR systems focus on formulating an adequate search query based on the conversational history. This query can then serve as the input to an existing, well-established retriever-ranker framework. We base our solution on CQR, leveraging its ability to integrate with existing ad-hoc search models, which has demonstrated significant practical value Dalton et al. (2022).
Earlier approaches to CQR attempted to select useful tokens from the conversation context Kumar and Callan (2020); Voskarides et al. (2020); Fang et al. (2022) or to train generative rewriter models with conversational sessions to mimic the human-rewritten query Yu et al. (2020); Lin et al. (2020); Vakulenko et al. (2021). To optimize query rewriting, some studies have adopted reinforcement learning Wu et al. (2022); Chen et al. (2022), or used the ranking signals with the rewriting model training Qian and Dou (2022); Mo et al. (2023a); Mao et al. (2023b). In addition, there have been endeavors to improve the conversion history quality through context denoising Lin et al. (2021b); Mao et al. (2022a); Krasakis et al. (2022); Mo et al. (2023b); Mao et al. (2023c), and data augmentation Dai et al. (2022); Mao et al. (2022b); Mo et al. (2024c). Unlike them, we enhance query rewriting by leveraging the NLP capabilities of open-source LLMs to reduce the ambiguity of conversational history. There have been multiple endeavors to integrate LLMs to solve traditional ad-hoc search sub-tasks Zhu et al. (2023), such as query expansion Wang et al. (2023); Gao et al. (2023), dense retrieval Ma et al. (2023); Wang et al. (2024b), and re-ranking Sun et al. (2023). About conversational search, Jin et al. (2023), Jang et al. (2023), and Chen et al. (2024a) attempt to improve CDR with unsupervised fine-tuning.  Mao et al. (2023a), Mo et al. (2024d), and Ye et al. (2023) explore how LLMs can understand users’ contextualized search intents via CQR. Unlike direct rewriting the query using LLMs, we investigate various approaches for integrating refined conversational history into CQR frameworks.
@@ 3.1 Task Formulation
Let =$\mathcal{H}=\{u_{k},r_{k}\}^{n}_{k=1}$ represent the user-system conversational history, where $u_{k}$ and $r_{k}$ are the user question and the system response at the $k$-th turn. Given a new user question $u_{n+1}$, the goal of a conversational search system is to return a set of passages $\mathcal{P}_{n+1}$ that are relevant to $\mathcal{H}$ and $u_{n+1}$, which would eventually help generate the model response $r_{n+1}$. To solve the main challenge of uncovering the real search intents hidden in the user’s context-dependent query, a conversational query rewriting (CQR) module has been commonly employed as an intermediate step to obtain a rewritten query $q_{n+1}$, which in turn is used as input to an off-the-shelf retriever. Recently, LLMs have become the default option for obtaining $q_{n+1}$ as follows:
@@ 3.2.1 Question Disambiguation
Users often expect a human-to-human level of interaction with modern conversational systems. They often use acronyms, ambiguous words, or coreference substitutes when asking questions, expecting native understanding and default reasoning capabilities from these systems. For a search system, the search intent is often unclear and ambiguous. Therefore, we propose a prompt to an LLM, denoted as $\mathcal{I}^{QD}$, that takes the conversational history $\mathcal{H}$ and the subsequent user question $u_{n+1}$ as input, to generate $u^{\prime}_{n+1}$, a self-contained and unambiguous version of $u_{n+1}$ which can substitute it in Eq. 1.
@@ 3.2.2 Response Expansion
In conversation sessions, it is common for model responses to be short and concise, especially for the factoid query. While brevity is often convenient for the user to acquire needed information, it makes the response less informative for a search system, which requires abundant rewrite/expansion resources. To handle this issue, we design a prompt $\mathcal{I}^{RE}$ which instructs an LLM to enrich the content of the last model response. The goal is to make it self-contained by leveraging the preceding conversational history. The enhanced history $\mathcal{H}^{\prime}$ is obtained by replacing the original response by $r^{\prime}_{n}$.
@@ 3.4 Search-Oriented Fine-tuning
Then, we select $q_{n+1}^{\prime}$, the one with the highest retrieval score $\mathcal{S}$ determined by an off-the-shelf retriever and relevance judgment from the set of pseudo-queries $Q^{\prime}_{n+1}$ as Eq 3, which is used as the supervision signal to fine-tune a rewriting model $\mathcal{M}(\mathcal{H}^{\prime}\oplus u_{n+1})=q_{n+1}^{\prime}$ by maximum likelihood estimation. It is important to mention that the process described in this section is conducted offline and performed once, for the purpose of generating pseudo-labeled queries to fine-tune a search-oriented query rewriter. During inference, $\mathcal{H}$ and $u_{n+1}$ serve as inputs for the fine-tuned model to generate the query $q_{n+1}$, and no calls are made to the LLM, so that the latency is not much affected.
@@ 4.2 Baselines
We compared our methods with a variety of systems that can mainly be classified into three categories. More precisely, we first compare against traditional systems that fine-tune small-scale CQR models (e.g., T5-base) including: QuReTeC Voskarides et al. (2020) T5QR Lin et al. (2020), CONQRR Wu et al. (2022), ConvGQR Mo et al. (2023a), EDIRCS Mao et al. (2023b). Then, we compare with the systems that fine-tune an LLM-based retriever, e.g., create the query and document representation by the ending token from a decoder-only model, including RepLLaMA Ma et al. (2023), E5-Mistral Wang et al. (2024b), and LLM-Embedder Zhang et al. (2023), or fine-tune an LLM-based CQR model as RETPO Yoon et al. (2024) and IterCQR Jang et al. (2023). Besides, we include the systems that directly obtain the rewritten query by prompting LLMs such as LLM-Aided IQR Ye et al. (2023), HyDE Gao et al. (2023), Query2doc Wang et al. (2023) and LLM4CS Mao et al. (2023a). Although not directly comparable, we report results of systems that fine-tune an ad-hoc search retriever for conversational scenarios, including the one without LLMs ConvDR Yu et al. (2021) and with LLMs InstructorR Jin et al. (2023). A detailed description of each aforementioned baseline is presented in Appendix B.3.
@@ 4.3 Implementation Details
We conduct experiments with the instruct-tuning variants [^2] Concretely, the version of LLMs we used are meta-llama/Llama-2-7b-chat-hf and mistralai/Mistral-7B-Instruct-v0.2 on https://huggingface.co/, respectively. of both LLaMA-2-7B Touvron et al. (2023) and Mistral-2-7B Jiang et al. (2023) as $\mathcal{LLM(.)}$ in Eq. 1 and Eq. 2. We experiment with both BM25 Robertson et al. (2009) sparse retriever and ANCE dense retriever Xiong et al. (2020). In addition, we use FlanT5-base [^3] We report our main results using FlanT5-base to ensure the results are comparable with previous studies. Chung et al. (2022) and large models as the backbone when fine-tuning a CQR model on TopiOCQA and QReCC. The fine-tuning process consists of 10 epochs with a learning rate of 1e-5 and a batch size of 8 for both datasets. More implementation details can be found in Appendix B.2.
@@ 6 Conclusion
Potential limitations of this work include not experimenting with larger open-source LLMs, such as the 56B Mixtral Mixtral AI team (2023) or 70B LLaM a, as well as other recent models like Gemma Team et al. (2024). Additionally, the study did not incorporate more closed-source models such as GPT-4 OpenAI (2023) or Claude AnthropicAI (2023) to further study the impact of history enhancement. This is mainly due to limitations in computation (open source) and financial (close source) resources. Despite the straightforward and significant gains, some design choices could be further analyzed to potentially boost the performance even more. For instance, adding a backoff filtering strategy could detect when the LLM is producing noisy outputs, or exploring approaches that interpolate between the use of human and pseudo-queries when its higher quality as training signals for CHIQ-FT. Besides, we have considered 5 directions of enhancement in this paper. More strategies can be incorporated so that other useful enhancements can be integrated.
@@ B.3 Baselines Details
E5-Mistral Wang et al. (2024b): A large ad-hoc retriever fine-tuned on top of Mistral-7B model on the synthetic dataset generated by ChatGPT-3.5 and MSMARCO.
@@ Appendix C Results of Mistral-2-7B
Table 7 and Table 8 show the performances of our methods using Mistral-2-7B as the backbone LLM, replacing LLama-2-7B as previously shown in Table 1 and Table 2, respectively. These results exhibit trends similar to those observed with LLama-2-7B, as discussed in Sections 5.1 and 5.2. More specifically, CHIQ-AD consistently outperforms CHIQ-FT across all settings, and combining the results of both methods (CHIQ-Fusion) yields the best performance. It is worth noting that results with Mistral-2-7B are systematically lower than the ones with LLama-2-7B on most datasets (except on CAsT-21). Besides, by comparing the fusion performance between Mistrial and LLaMa, we notice that in most cases when the gap between CHIQ-AD and CHIQ-FT is large, CHIQ-Fusion results are either slightly better or worse than CHIQ-AD. This is mainly because the poor quality of the rank list obtained by CHIQ-FT negatively impacts the one from CHIQ-AD. However, when the gap is smaller, we notice a significant gain for CHIQ-Fusion, suggesting that both variants are generating good and complementary rank lists. It will be interesting to investigate how we can better take advantage of CHIQ-AD and CHIQ-FT in an adaptive fusion. Nevertheless, performing QR on top of the enhanced history with our approach still outperforms most other settings and datasets.